# High-dimensional inference of radio interferometer beam patterns I: Parametric model of the HERA beams

### v1

@@ 5 Summary and Conclusion
When we perform this simulation, we find that there is no sign of deleterious spurious spectral structure. For the Vivaldi feed, performance at high delays greater than 2000 ns is nearly identical. Performance at high delays for the Phase I dipole is slightly worse than the CST simulation, but still within the dynamic range requirements set by the reionization signal. We observed the largest differences for both feed types at intermediate delays in between the Nyquist delay of the CST simulation (500 ns) and high delays. The simulated beams exhibit stronger power at these delays, which we suppose is likely to be an interpolation artefact since structure at delays beyond the Nyquist delay should not be recovered by a spline interpolation except by coincidence.

### v2

@@ 5 Summary and Conclusion
When we perform this simulation, we find that there is no sign of deleterious spurious spectral structure. For both feeds, performance at high delays greater than 2000 ns is nearly identical to the simulated beams. Performance at intermediate to high delays is within the dynamic range requirements set by the reionization signal. We observed the largest differences for both feed types at intermediate delays surrounding the Nyquist delay of the CST simulation (500 ns) and around 1500 ns. The simulated beams exhibit stronger power at these delays. We suppose this is likely to be an interpolation artefact, since structure at delays beyond the Nyquist delay should not be recovered by a spline interpolation except by coincidence, and various simulations suggest the structure is not an intrinsic structure of the simulated beams or sky model.